Let A and B are the bivariate continuous random variables. The joint probability density function is given by 𝑓(π‘Ž, 𝑏) = { π‘Ž 2𝑏π‘₯ 0 < π‘Ž < 3, 0 < 𝑏 < 2; 0 otherwise. (i) Determine the value of x for which the function can serve as a joint probability density function. (ii) Find the marginal density function of B. (iii) Find 𝑓(π‘Ž < 1|𝑏 = 1) Β 


1
Expert's answer
2021-07-19T05:41:23-0400

(i)


"\displaystyle\int_{-\infin}^{\infin}\displaystyle\int_{-\infin}^{\infin}f(a,b)dbda=\displaystyle\int_{0}^{3}\displaystyle\int_{0}^{2}a^2bxdbda"

"=x\displaystyle\int_{0}^{3}a^2[\dfrac{b^2}{2}]\begin{matrix}
 2 \\
 0
\end{matrix}da=2x\displaystyle\int_{0}^{3}a^2da"

"=2x[\dfrac{a^3}{3}]\begin{matrix}
 3 \\
 0
\end{matrix}=18x=1=>x=\dfrac{1}{18}"


"f(a,b) = \begin{cases}
 \dfrac{1}{18}a^2b, &0\leq a\leq 3, 0\leq b\leq 2 \\
 0, & otherwise
\end{cases}"

(ii)


"f_B(b)=\displaystyle\int_{-\infin}^{\infin}f(a,b)da=\displaystyle\int_{0}^{3} \dfrac{1}{18}a^2bda"

"= \dfrac{1}{18}b[\dfrac{a^3}{3}]\begin{matrix}
 3 \\
 0
\end{matrix}=\dfrac{1}{2}b"

"f_B(b) = \begin{cases}
 \dfrac{1}{2}b, & 0\leq b\leq 2 \\
 0, & otherwise
\end{cases}"

(iii) For "0\leq b\leq 2" we obtain


"f(a|b)=\dfrac{f(a, b)}{f_B(b)}=\dfrac{1}{9}a^2,0\leq a\leq 3"

"f(a<1|b=1)=\displaystyle\int_{0}^{1} \dfrac{1}{9}a^2da=\dfrac{1}{27}"